OpenStudy (anonymous):

Part 1: Solve the following equation using one of the methods listed below. Part 2: Using complete sentences, explain why you chose the method you used. Equation: 0 = x2 + 10x + 5

OpenStudy (anonymous):

Methods: Completing the Square Factoring Quadratic Formula Graphing

OpenStudy (anonymous):

@JuniorEinstein1

OpenStudy (anonymous):

@jcpd910

OpenStudy (anonymous):

0 = x2 + 10x + 5 Let's use the quadratic formula, it's the most fun.

OpenStudy (anonymous):

\[x = \frac{ -b \sqrt{b^{2} - 4ac} }{ 2a }\]

OpenStudy (anonymous):

0 = x2 + 10x + 5 0 = ax^2 + bx + c a = 1 b = 10 c = 5

OpenStudy (anonymous):

ok i got x=[ - 5 ± 2√5] / 2

OpenStudy (anonymous):

is it right

OpenStudy (anonymous):

\[x = \frac{ -10 \sqrt{10^{2} - 4(1)(5)} }{ 2(1) }\]

OpenStudy (anonymous):

and what should i write for part 2

OpenStudy (anonymous):

Part 2: Using complete sentences, explain why you chose the method you used. I used it because I am more familiar with it than the other methods.

OpenStudy (anonymous):

Now I'ma solve for x

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

i think i am right on my answer

OpenStudy (anonymous):

i gotta go bye

OpenStudy (anonymous):

\[x = \frac{ -10 \sqrt{10^{2} - 4(1)(5)} }{ 2(1) }\] \[x = \frac{ -10 \sqrt{100 -20} }{ 2 }\] \[x = \frac{ -10 \sqrt{80} }{ 2 }\] \[x = -20\sqrt{5}\]